OpenStudy (anonymous):

The perimeter of a rectangle is 136inches. The length exceeds the width by 56 inches. Find the length and width

OpenStudy (anonymous):

let width = x in then length = x + 56 in Perimeter = 2L + 2B 136 = 2(x + 56) + 2(x) 136 = 2x + 112 + 2x 136 - 112 = 4x 24/4 = x x = 6 so width is = 6 in length is 6 + 56 = 62 in

OpenStudy (anonymous):

if u r satisfied with the answer pls click GOOD ANSWER button for me ( :))